OpenStudy (anonymous):

how to solve 3b^2+9b=0

OpenStudy (anonymous):

I got -3 as an answer. is that correct?

OpenStudy (anonymous):

Factor \[3b^2+9b=0\]

OpenStudy (anonymous):

-3 is a solution. There is another solution.

OpenStudy (anonymous):

\[3b^2+9b=0\] \[3b(b+3)=0\]

OpenStudy (anonymous):

0,-3 3b^2+9b= 0 b(3b+9)=0 b=0 0=3b+9 -9=3b -9/3 -3=b

OpenStudy (anonymous):

\[3b = 0\] \[\frac{ 3b }{ 3 } =\frac{ 0 }{ 3 }\] \[b = 0\]

OpenStudy (anonymous):

\[b + 3 = 0\] \[b = -3\]

OpenStudy (anonymous):

@Luis_Rivera plug -3 in, and note the +9b. It works. There's also a solution if you factor out 3b=0: b can equal zero as well.

OpenStudy (anonymous):

\[\sqrt{?} \] square root

mathslover (mathslover):

@Luis_Rivera "You can not have a negative square root" :o , \(\sqrt{-1} = iota\)

mathslover (mathslover):

Though, it is imaginary but we can have negative square roots.

OpenStudy (anonymous):

That was hardly a mathematical concept. I'm not clarifying anything for you, I'm clarifying for the person who asked the question. Try not to let making slight errors blind you to reality, and please try not to feel offended when I answer questions correctly.

OpenStudy (anonymous):

And you try to get used to the idea that i'll respond to any comment i feel the need to. If you dont like it, then keep your self-righteousness to yourself. and dont send me silly and childish private messages. And by the way, don't bother replying to this. I may give importance to your response when you reach my level. Right now you're just another user.

OpenStudy (anonymous):

I find public bickering distasteful. Also, I would like to point out that my statistics are better than yours when you normalize them to the days we've been online, so by that measure I'm at your level.

OpenStudy (anonymous):

Oh, and I apologize to anybody who is being spammed with notifications because of this.

OpenStudy (anonymous):

what on earth ?